### v1

@@ Abstract
footnotetext: This work has been submitted to Nature Electronics for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
@@ 1.1 The Adiabatic Neuron: The Computational Workhorse
Figure 2: The ACN computational workhorse: a) Schematic of a DTSC ACN with a single power clock signal $PC$, an NN-bit input vector $x$, NN synapse capacitors, threshold logic unit and ACN output, $y$; b) Experimentally measured oscilloscope trace of the primary ACN signals illustrating operational timings of one of the 12 1-bit input ACNs in our chip; c) 130nmnm CMOS layout of the same ACN.
@@
A prominent feature of the ACN is its dual switched-capacitor arrays, or trees, representing the positive and negative-weighted synapses, respectively. Each of the $N$ synapse capacitors is connected to the sinusoidal PC dependent on a switch controlled by the corresponding bit of the $N$-bit ACN input, $x$; otherwise, the capacitor remains grounded. The membrane nodes, $v_{m}^{\pm}$, then perform multi-input MAC operations via the principle of capacitive charge redistribution and voltage division. Bias capacitors, ballast capacitors and the connections to $VBN$ and $VBP$ DC voltages enable the implementation of neuron biases, $v_{m}^{\pm}$ node swing moderation and periodic $v_{m}^{\pm}$ node resetting, respectively. Detailed explanations can be found in [Smart2025].
@@ 1.2 An ACN Network
Fig. 3b shows the architectural block diagram of our complete ACNN chip. The 64-bit test images are clocked in via a 1MHzMHz SPI serial interface to the ACNN Deserializer shift register. The resulting 64-bit wide parallel input signal is then provided to the first layer of 12 mapped and independently instantiated ACNs. The outputs generated by the first layer are routed to a second layer of 4 ACNs through an intermediate routing layer. The 4-bit inference output, $QL2$, is subsequently transmitted to the external chip pads for off-chip readout. The 12-bit outputs from our first layer are also routed to external chip pads as $QL1$, for debugging purposes. The two ACN layers represent about 51pFpF of distributed capacitance, driven by an externally provided single power clock, $PC1$, via the control logic, which also provides the two associated non-overlapping TL clock signals, $CCLK1$ and $CCLK2$. A description and purpose of the other control logic signals in our chip are provided in Method 3.3.
@@ 1.3.2 Cross-Chip Energy Efficiency
Total ACNN Energy ($pJ$)
@@
Total CCNN Energy ($pJ$)
@@ 3.3 Internal configuration and control of the ACNN pipeline
During testing each 64-bit (8x8 1-bit pixel) test image was clocked in serially at $1MHz$ to the ACNN chip from an external SPI master, using standard SPI data signal ($DATA$), serial clock ($SCLK$) and chip select ($CS$). The active-low reset signal, $RST$, is a proprietary signal used to reset the $Deserializer$ logic ready for a new test image. The $Deserializer$ outputs a 64-bit parallel data bus, which acts as control signals to the 64 adiabatic switches in each first-layer ACN.
Finally, analog DC voltages $VBP$ and $VBN$ were routed from external pads to every ACN. During a reset period, when the signal $START=0$, Single Pole Single Throw (SPST) switches in each ACN are closed, allowing $VBP/VBN$ to be applied to the $v_{m}^{\pm}$ nodes. After $START$ goes high, these SPST switches are reopened, and computation can begin. The 12-bit output from the first layer, $QL1$, becomes valid after the first $PC1$ cycle. These bits then feed into the routing layer and, subsequently, layer 2, as well as being connected to external pads for debugging. The final computation, the image classification, becomes valid on $QL2$ after the second $PC1$ cycle, using a 4-bit one-hot encoding scheme.
@@ 3.4 PCB PC clock generation design and operation
The Power Clock (PC) for our ACNN chip is generated on the PCB whose circuit is shown simplified in Extended Data Fig. 6a. The circuit consists of a tank capacitor $C_{tank}=100nF$, inductor, $L_{pc}=0.39mH$, $C_{pc}=25pF$ and DC supply voltage, $V_{sup}=0.62V$. The switch $CHRG$ is used to connect the DC supply voltage to the tank capacitor, facilitating its charging phase. The tank capacitor can be charged at the beginning of each operational cycle by closing the switch (CHRGCHRG=logic high). Once the tank capacitor is charged the switch is opened to isolate the supply, thereby completing the charging sequence and preparing the system for the subsequent operational phase.
Following the charging phase, switch $PULSE$ is closed, connecting the tank capacitor to the inductor. This couples the tank capacitor with the inductor, initiating an LC resonance that allows energy exchange between the two components. This generates the required sinusoidal voltage at node $V_{pc}$. During this phase, switch $PULSEb$ remains open to avoid disrupting the oscillation. The duration for which switch $PULSE$ remains closed governs the frequency of the generated sinusoidal waveform. Once the oscillation completes, switch $PULSE$ is opened and switch $PULSEb$ is subsequently closed to reset the $V_{pc}$ node to ground, establishing the ideal phase spacing between successive sinusoidal cycles. The complete sequence of tank capacitor charging and the timed control of switches $PULSE$ and $PULSEb$ is illustrated in the simplified timing diagram shown in Extended Data Fig. 6b.
Real PCG implementations, as implemented in our daughter board, include non-idealities, such as unwanted overshoot, reduced PC peaks, glitches and delays in TL clock generation signals, but also more subtlety, potential changes in frequency with capacitive load variations, as shown in Extended Data Fig. 6c. A reduced maximum peak voltage of 1.47VV in $PC1$ was coupled with a 150mVmV of unwanted undershoot. Small glitches and delays in the $CCLK$ generation derived from the PC are also possible.
@@ 3.5 Chip test configuration and custom ACNN daughter board
The ArC2 interfaces with a custom daughter board that hosts the DUT and generates the $PC$ signals using two LC ($L_{pc}=0.390mH$ and $C_{pc}=25pF$) oscillator circuits, as required by our ACNN chip. To replace lost charge during computation, each LC circuit was topped up from a charged $C_{tank}=100nF$ tank capacitor, acting effectively as a battery. The board also generates the associated $CCLK$ TL sampling signals for each PC, as previously discussed.
The timing diagram in Extended Data Fig. 7d shows a single computational operation. The ArC2 first generates a $10\mu s$ CHRGCHRG strobe to charge both tank capacitors, up to the supply voltage, of around 0.62VV. This tank capacitor can be topped up every operation, or just once, depending on the type of test. After raising the $RST$ signal, the 64-bit test arrows8 image is injected serially into the DUT via an SPI connection. Once loaded, the daughter board is instructed by the ArC2 to start generating the PCPC/CCLKCCLK signals, 200nsns after the rising edge of $TRG$. A fixed number of sinusoidal-like peaks is then generated on each of the PCs: three on $PC1$ and four on $PC2$. After $START$ is raised, and the computation is completed, software reads the 4-bit $QL2$ classification output.
@@ 3.6 Measuring energy dissipation in the ACNN chip
A knock-on effect of draining the tank capacitors was that, as each experiment progressed, the $PC$ signals were subject to decay in their peak amplitudes over time. This meant the computations were increasingly susceptible to miscalculation. In this paper, the dissipated energy per synapse operation, $E_{SOP}$, was determined by
@@ Appendix A Extended Data
Figure 7: ACNN chip test setup: a) ArC2 and ACNN daughter board block diagram controlled via a laptop and two USB connections. Laptop configures the ArC2 to provide VBNVBN and $VBP$, typically 0VV, analog voltages, accurately schedules the digital IO signals for PC clock generation circuits and configures/controls ACNN operations. The 64-bit test patterns are written to the ACNN deserializer via an FT232H serial chip; b) Photograph of the actual setup; c) A prototype ACNN test chip; d) Test system timing diagram for a single image classification operation. The PCB tank capacitor is charged first, followed by the loading of the 64-bit test data, before both the PC clock generators and ACNN are initiated by the TRGTRG and STARTSTART signals, respectively. After computation, the ACNN L2 classification outputs are read back to the software via the FT232H chip.
@@
footnotetext: This work has been submitted to Nature Electronics for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.

### v2

@@ Abstract
footnotetext: This work has been submitted to Nature Communications for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
@@ 1.1 The Adiabatic Neuron: The Computational Workhorse
Figure 2: The ACN computational workhorse: a) Schematic of a DTSC ACN with a single power clock signal PCPC, an NN-bit input vector $x$, NN synapse capacitors, threshold logic unit and ACN output, $y$; b) Experimentally measured oscilloscope trace of the primary ACN signals illustrating operational timings of one of the 12 1-bit input ACNs in our chip; c) 130nmnm CMOS layout of the same ACN.
@@
A prominent feature of the ACN is its dual switched-capacitor arrays, or trees, representing the positive and negative-weighted synapses, respectively. Each of the NN synapse capacitors is connected to the sinusoidal PC dependent on a switch controlled by the corresponding bit of the NN-bit ACN input, $x$; otherwise, the capacitor remains grounded. The membrane nodes, $v_{m}^{\pm}$, then perform multi-input MAC operations via the principle of capacitive charge redistribution and voltage division. Bias capacitors, ballast capacitors and the connections to VBNVBN and VBPVBP DC voltages enable the implementation of neuron biases, $v_{m}^{\pm}$ node swing moderation and periodic $v_{m}^{\pm}$ node resetting, respectively. Detailed explanations can be found in [Smart2025].
@@ 1.2 An ACN Network
Fig. 3b shows the architectural block diagram of our complete ACNN chip. The 64-bit test images are clocked in via a 1MHzMHz SPI serial interface to the ACNN Deserializer shift register. The resulting 64-bit wide parallel input signal is then provided to the first layer of 12 mapped and independently instantiated ACNs. The outputs generated by the first layer are routed to a second layer of 4 ACNs through an intermediate routing layer. The 4-bit inference output, $QL2$, is subsequently transmitted to the external chip pads for off-chip readout. The 12-bit outputs from our first layer are also routed to external chip pads as $QL1$, for debugging purposes. The two ACN layers represent about 51pFpF of distributed capacitance, driven by an externally provided single power clock, $PC1$, via the control logic, which also provides the two associated non-overlapping TL clock signals, $CCLK1$ and $CCLK2$. A description and purpose of the other control logic signals in our chip are provided in Method $3.$
@@ 1.3.2 Cross-Chip Energy Efficiency
Total ACNN Energy (pJpJ)
@@
Total CCNN Energy (pJpJ)
@@ 3.3 Internal configuration and control of the ACNN pipeline
During testing each 64-bit (8x8 1-bit pixel) test image was clocked in serially at $1MHz$ to the ACNN chip from an external SPI master, using standard SPI data signal (DATADATA), serial clock (SCLKSCLK) and chip select (CSCS). The active-low reset signal, RSTRST, is a proprietary signal used to reset the DeserializerDeserializer logic ready for a new test image. The DeserializerDeserializer outputs a 64-bit parallel data bus, which acts as control signals to the 64 adiabatic switches in each first-layer ACN.
Finally, analog DC voltages VBPVBP and VBNVBN were routed from external pads to every ACN. During a reset period, when the signal $START=0$, Single Pole Single Throw (SPST) switches in each ACN are closed, allowing $VBP/VBN$ to be applied to the $v_{m}^{\pm}$ nodes. After STARTSTART goes high, these SPST switches are reopened, and computation can begin. The 12-bit output from the first layer, $QL1$, becomes valid after the first $PC1$ cycle. These bits then feed into the routing layer and, subsequently, layer 2, as well as being connected to external pads for debugging. The final computation, the image classification, becomes valid on $QL2$ after the second $PC1$ cycle, using a 4-bit one-hot encoding scheme.
@@ 3.4 PCB PC clock generation design and operation
The Power Clock (PC) for our ACNN chip is generated on the PCB whose circuit is shown simplified in Extended Data Fig. 6a. The circuit consists of a tank capacitor $C_{tank}=100nF$, inductor, $L_{pc}=0.39mH$, $C_{pc}=25pF$ and DC supply voltage, $V_{sup}=0.62V$. The switch CHRGCHRG is used to connect the DC supply voltage to the tank capacitor, facilitating its charging phase. The tank capacitor can be charged at the beginning of each operational cycle by closing the switch (CHRGCHRG=logic high). Once the tank capacitor is charged the switch is opened to isolate the supply, thereby completing the charging sequence and preparing the system for the subsequent operational phase.
Following the charging phase, switch PULSEPULSE is closed, connecting the tank capacitor to the inductor. This couples the tank capacitor with the inductor, initiating an LC resonance that allows energy exchange between the two components. This generates the required sinusoidal voltage at node $V_{pc}$. During this phase, switch PULSEbPULSEb remains open to avoid disrupting the oscillation. The duration for which switch PULSEPULSE remains closed governs the frequency of the generated sinusoidal waveform. Once the oscillation completes, switch PULSEPULSE is opened and switch PULSEbPULSEb is subsequently closed to reset the $V_{pc}$ node to ground, establishing the ideal phase spacing between successive sinusoidal cycles. The complete sequence of tank capacitor charging and the timed control of switches PULSEPULSE and PULSEbPULSEb is illustrated in the simplified timing diagram shown in Extended Data Fig. 6b.
Real PCG implementations, as implemented in our daughter board, include non-idealities, such as unwanted overshoot, reduced PC peaks, glitches and delays in TL clock generation signals, but also more subtlety, potential changes in frequency with capacitive load variations, as shown in Extended Data Fig. 6c. A reduced maximum peak voltage of 1.47VV in $PC1$ was coupled with a 150mVmV of unwanted undershoot. Small glitches and delays in the CCLKCCLK generation derived from the PC are also possible.
@@ 3.5 Chip test configuration and custom ACNN daughter board
The ArC2 interfaces with a custom daughter board that hosts the DUT and generates the PCPC signals using two LC ($L_{pc}=0.390mH$ and $C_{pc}=25pF$) oscillator circuits, as required by our ACNN chip. To replace lost charge during computation, each LC circuit was topped up from a charged $C_{tank}=100nF$ tank capacitor, acting effectively as a battery. The board also generates the associated CCLKCCLK TL sampling signals for each PC, as previously discussed.
The timing diagram in Extended Data Fig. 7d shows a single computational operation. The ArC2 first generates a $10\mu s$ CHRGCHRG strobe to charge both tank capacitors, up to the supply voltage, of around 0.62VV. This tank capacitor can be topped up every operation, or just once, depending on the type of test. After raising the RSTRST signal, the 64-bit test arrows8 image is injected serially into the DUT via an SPI connection. Once loaded, the daughter board is instructed by the ArC2 to start generating the PCPC/CCLKCCLK signals, 200nsns after the rising edge of TRGTRG. A fixed number of sinusoidal-like peaks is then generated on each of the PCs: three on $PC1$ and four on $PC2$. After STARTSTART is raised, and the computation is completed, software reads the 4-bit $QL2$ classification output.
@@ 3.6 Measuring energy dissipation in the ACNN chip
A knock-on effect of draining the tank capacitors was that, as each experiment progressed, the PCPC signals were subject to decay in their peak amplitudes over time. This meant the computations were increasingly susceptible to miscalculation. In this paper, the dissipated energy per synapse operation, $E_{SOP}$, was determined by
@@ Appendix A Extended Data
Figure 7: ACNN chip test setup: a) ArC2 and ACNN daughter board block diagram controlled via a laptop and two USB connections. Laptop configures the ArC2 to provide VBNVBN and VBPVBP, typically 0VV, analog voltages, accurately schedules the digital IO signals for PC clock generation circuits and configures/controls ACNN operations. The 64-bit test patterns are written to the ACNN deserializer via an FT232H serial chip; b) Photograph of the actual setup; c) A prototype ACNN test chip; d) Test system timing diagram for a single image classification operation. The PCB tank capacitor is charged first, followed by the loading of the 64-bit test data, before both the PC clock generators and ACNN are initiated by the TRGTRG and STARTSTART signals, respectively. After computation, the ACNN L2 classification outputs are read back to the software via the FT232H chip.